# s1: Simple test-time scaling

### v1

@@ 1 Introduction
Performance improvements of language models (LMs) over the past years have largely relied on scaling up train-time compute using large-scale self-supervised pretraining (Kaplan et al., 2020; Hoffmann et al., 2022). The creation of these powerful models has set the stage for a new scaling paradigm built on top of them: test-time scaling. The aim of this approach is to increase the compute at test time to get better results. There has been much work exploring this idea (Snell et al., 2024; Wu et al., 2024b; Welleck et al., 2024), and the viability of this paradigm was recently validated by OpenAI o1 (OpenAI, 2024). o1 has demonstrated strong reasoning performance with consistent gains from scaling test-time compute. OpenAI describes their approach as using large-scale reinforcement learning (RL) implying the use of sizable amounts of data (OpenAI, 2024). This has led to various attempts to replicate their models relying on techniques like Monte Carlo Tree Search (Gao et al., 2024b; Zhang et al., 2024a), multi-agent approaches (Huang et al., 2024b), and others (Huang et al., 2024b; Gao et al., 2024b; Zhang et al., 2024a; Wang et al., 2024a). Among these approaches, DeepSeek R1 (DeepSeek-AI et al., 2025) has successfully replicated o1-level performance, also employing reinforcement learning via millions of samples and multiple training stages. However, despite the large number of o1 replication attempts, none have openly replicated a clear test-time scaling behavior. Thus, we ask: what is the simplest approach to achieve both test-time scaling and strong reasoning performance?
@@
In summary, our contributions are: We develop simple methods for creating a sample-efficient reasoning dataset (§2) and test-time scaling (§3); Based on these we build s1-32B which is competitive with o1-preview (§4); We ablate subtleties of data (§5.1) and test-time scaling (§5.2). We end with a discussion to motivate future work on simple reasoning (§6). Our code, model, and data will be open-source.
@@ 3.2 Metrics
Figure 4: Sequential and parallel test-time scaling. (a): Budget forcing shows clear scaling trends and extrapolates to some extent. For the three rightmost dots, we prevent the model from stopping its thinking 2/4/6 times, each time appending “Wait” to its current reasoning trace. (b): For Qwen2.5-32B-Instruct we perform 64 evaluations for each sample and visualize the performance when majority voting across 2, 4, 8, 16, 32, and 64 of these.
@@ 4.1 Setup
We benchmark s1-32B against: OpenAI o1 series (OpenAI, 2024), which are closed-source models that popularized the idea of test-time scaling; DeepSeek r1 series (Team, 2024a), which are open-weight reasoning models with up to o1-level performance, concurrently released to ours; Qwen’s QwQ-32B-preview (Team, 2024b), a 32B open-weight reasoning model without disclosed methodology; Sky-T1-32B-Preview, which is an open-weight and open reasoning data model distilled from QwQ-32B-preview; Google Gemini 2.0 Flash Thinking Experimental, the API that we are distilling from. As there are no official evaluation scores, we use the Gemini API to benchmark it ourselves. However, the “recitation error” associated with the Gemini API makes evaluation challenging. [^2] https://github.com/google/generative-ai-docs/issues/257 We circumvent this, by manually inserting all 30 AIME24 questions in its web interface where the error does not appear. However, we leave out MATH500 (500 questions) and GPQA Diamond (198 questions), thus they are N.A. in LABEL:tab:perf. Our model, s1-32B, is fully open including weights, reasoning data, and code.
@@
Model # ex. AIME 2024 MATH- 500 GPQA Diamond
@@ 4.2 Performance
In Figure 2 (right) and LABEL:tab:perf we compare s1-32B with other models. We find that s1-32B is the most sample-efficient open data reasoning model. It performs significantly better than our base model (Qwen2.5-32B-Instruct) despite just training it on an additional 1,000 samples. The concurrently released r1-32B shows stronger performance than s1-32B while also only using SFT (DeepSeek-AI et al., 2025). However, it is trained on 800 $\times$ more reasoning samples. It is an open question whether one can achieve their performance with just 1,000 samples. Finally, our model nearly matches Gemini 2.0 Thinking on AIME24, as s1-32B is distilled from Gemini 2.0, this shows that our distillation procedure was likely effective.
@@
[AIME24] Alice and Bob play the following game. A stack of $n$ tokens lies before them. The players take turns with Alice going first. On each turn, the player removes either $1$ token or $4$ tokens from the stack. Whoever removes the last token wins. Find the number of positive integers $n$ less than or equal to $2024$ for which there exists a strategy for Bob that guarantees that Bob will win the game regardless of Alice’s play. The problem asks for the number of positive integers $n\leq 2024$ such that Bob has a winning strategy. This is a typical impartial game that can be analyzed using the Sprague-Grundy theorem or by directly finding the winning and losing positions. Let $W$ be the set of winning positions (number of tokens $n$ such that the player whose turn it is can force a win) and $L$ be the set of losing positions. […] Let $P(n)$ be the statement ”the player whose turn it is when there are $n$ tokens has a winning strategy. $P(0)$ is false (the player whose turn it is and faces 00 tokens has already lost). Consider the small values of $n$: $n=1$: The player whose turn it is can take $1$ token and win. So $P(1)$ is true. $1\in W$. $n=2$: The player whose turn it is can take $1$ token, leaving $1$ token for the opponent. Since $1\in W$, the opponent wins. So $P(2)$ is false. $2\in L$. $n=3$: The player whose turn it is can take $1$ token, leaving $2$ tokens for the opponent. Since $2\in L$, the player whose turn it is wins. So $P(3)$ is true. $3\in W$.[…] We are looking for the number of positive integers $n\leq 2024$ for which Bob has a winning strategy. This means that if the game starts with $n$ tokens, and Alice goes first, Bob can win. This happens if and only if $n$ is a losing position (because if $n\in L$, then Alice’s first move must lead to a winning position for Bob). […] Let’s try to find a pattern in the losing positions $L=\{0,2,5,7,10,12,15,17,20,22,...\}$. Differences between consecutive terms: $2,3,2,3,2,3,2,3,2,...$ The pattern of differences is alternating $2,3$. […] The losing positions are of the form $5m$ or $25m+$ 2, where $m\geq 0$. We need to find the number of positive integers $n\leq 2024$ that are in the set of losing positions $L$. […] Case 1: $n=5m$, where $m\geq 1$. $1\leq 5m\leq 2024$ $404.81/5\leq m\leq 2024/5=$ 404.8 Since $m$ is an integer, $1\leq m\leq 404$. The number of values of $m$ is $404$. Case 2: $n=5m+2$, where $m\geq 0$. $5m+2\leq 2024$ $5m\leq 2022$ $m\leq 2022/5=404.4$ Since $m$ is an integer, $0\leq m\leq 404$. The number of values of $m$ is $404-0+1=405$. The total number of positive integers $n\leq 2024$ that are losing positions is $404+405=\boxed{809}$. [MATH] Find the projection of $a$ onto $b=\begin{pmatrix}2\\ 6\\ 3\end{pmatrix}$ if $a\cdot b=8.$ The projection of vector $\mathbf{a}$ onto vector $\mathbf{b}$ is given by the formula: proj_b a = abb2 b We are given $\mathbf{b}=\begin{pmatrix}2\\ 6\\ 3\end{pmatrix}$ and $\mathbf{a}\cdot\mathbf{b}=8.$ First, we need to find the magnitude squared of $\mathbf{b}$, which is $\|\mathbf{b}\|^{2}$. ∥b∥^2 = bb = 2^2 + 6^2 + 3^2 = 4 + 36 + 9 = 49. Now we can plug the values into the projection formula: proj_b a = 849 b Substitute the vector $\mathbf{b}$: proj_b a = 849 (2 6 3 ) Multiply the scalar by each component of the vector: proj_b a = (849×2 849×6 849×3 ) = (164948492449) Final Answer: The final answer is ($164948492449$ ) [GPQA] A quantum mechanical particle of mass $m$ moves in two dimensions in the following potential, as a function of the polar coordinates $(r,\theta)$: V(r, θ) = 12 k r^2 + 32 k r^2 cos^2(θ) Find the energy spectrum. Hint: Write the potential in Cartesian coordinates. The problem asks for the energy spectrum of a particle in a given 2D potential. The potential is given in polar coordinates $(r,\theta)$: V(r, θ) = 12 kr^2 + 32 kr^2 cos^2(θ). We can rewrite the potential as: V(r, θ) = 12 kr^2 (1 + 3cos^2(θ)). The Hamiltonian for a particle in 2D is given by: H = -ℏ22m ( ∂2∂x2 + ∂2∂y2 ) + V(x, y). In polar coordinates $(r,\theta)$, the Laplacian operator is: ∇^2 = 1r ∂∂r (r ∂∂r ) + 1r22θ2. So the Hamiltonian in polar coordinates is: H = -ℏ22m [ 1r ∂∂r (r ∂∂r ) + 1r22θ2 ] + V(r, θ). The time-independent Schrödinger equation is $H\psi=E\psi$. We can try to separate variables by assuming a solution of the form $\psi(r,\theta)=R(r)\Theta(\theta)$. […] This confirms the energy spectrum is: E = ℏkm (2n_x + n_y + 32). Final Answer: $E=\hbar\sqrt{\frac{k}{m}}\left(2n_{x}+n_{y}+\frac{3}{2}\right)$.
@@
Figure 5: Example model outputs. We pick a question from AIME24 (left), MATH (middle), and GPQA (right), where our model generates the correct answer. The black text is the prompt, the light blue text is the reasoning trace, and the blue text is the answer of s1-32B. The gray ellipsis […] indicates that the text was trimmed to fit this page, but the generated text is actually longer.
@@ 5.1 Data Quantity, Diversity, and Difficulty
Table 2: s1K data ablations. We budget force (BF) a maximum of 32,000 thinking tokens for scores in this table. This performs slightly better than the scores without BF (LABEL:tab:perf) as it allows the model to finish with a best guess when stuck in an infinite loop. We report 95% Wilson confidence intervals with ±.
Model AIME 2024 MATH- 500 GPQA Diamond
1K-random 36.7 ± 14.8 90.6 ± 2.3 52.0 ± 6.9
1K-diverse 26.7 ± 12.5 91.2 ± 2.2 54.6 ± 6.8
1K-longest 33.3 ± 14.0 90.4 ± 2.3 59.6 ± 6.6
59K-full 53.3 ± 16.8 92.8 ± 2.0 58.1 ± 6.9
s1K 50.0 ± 16.8 93.0 ± 1.9 57.6 ± 6.7
@@ 5.2 Test-time scaling methods
Model AIME 2024 MATH -500 GPQA Diamond
@@ 6.1 Sample-efficient reasoning
To evaluate and push the limits of these models, increasingly challenging benchmarks have been introduced, such as Olympiad-level science competitions (He et al., 2024a; Jain et al., 2024a; Zhong et al., 2023) and others (Srivastava et al., 2023; Glazer et al., 2024; Su et al., 2024; Kim et al., 2024; Phan et al., 2025). To enhance models’ performance on reasoning-related tasks, researchers have pursued several strategies: Prior works have explored continuing training language models on specialized corpora related to mathematics and science (Azerbayev et al., 2023; Yang et al., 2024), sometimes even synthetically generated data (Yu et al., 2024). Others have developed training methodologies specifically aimed at reasoning performance (Zelikman et al., 2022, 2024; Luo et al., 2025; Yuan et al., 2025). Another significant line of work focuses on prompting-based methods to elicit and improve reasoning abilities, including methods like Chain-of-Thought prompting (Wei et al., 2023; Yao et al., 2023a, b; Bi et al., 2023; Fu et al., 2023; Zhang et al., 2024b; Hu et al., 2024). These combined efforts aim to advance the reasoning ability of language models, enabling them to handle more complex and abstract tasks effectively.
@@ Methods
As we introduce in §3, we differentiate two methods to scale test-time compute: parallel and sequential. The former relies on multiple solution attempts generated in parallel and selecting the best outcome via specific criteria. These criteria include choosing the most frequent response for majority voting or the best response based on an external reward for Best-of-N (Brown et al., 2024; Irvine et al., 2023; Snell et al., 2024). Unlike repeated sampling, previous sequential scaling methods let the model generate solution attempts sequentially based on previous attempts, allowing it to refine each attempt based on previous outcomes (Snell et al., 2024; Hou et al., 2025; Lee et al., 2025). Tree-based search methods (Gandhi et al., 2024; Wu et al., 2024a) offer a hybrid approach between sequential and parallel scaling, such as Monte-Carlo Tree Search (MCTS) (Liu et al., 2024; Zhang et al., 2023; Zhou et al., 2024; Choi et al., 2023) and guided beam search (Xie et al., 2023). REBASE (Wu et al., 2024a) employs a process reward model to balance exploitation and pruning during tree search. Empirically, REBASE has been shown to outperform sampling-based methods and MCTS (Wu et al., 2024a). Reward models (Lightman et al., 2023; Wang et al., 2024b, c) play a key role in these methods. They come in two variants: outcome reward models and process reward models. Outcome reward models (Xin et al., 2024; Ankner et al., 2024) assign a score to complete solutions and are particularly useful in Best-of-N selection, while process reward models (Lightman et al., 2023; Wang et al., 2024b; Wu et al., 2024a) assess individual reasoning steps and are effective in guiding tree-based search methods.
@@
Parallel scaling offers one solution to the limits of sequential scaling, thus we augment our sequentially scaled model with two methods: (I) Majority voting: After generating $k$ solutions, the final solution is the most frequent one across generations; (II) Tree search via REBASE: We use the REBASE process reward model, which is initialized from LLaMA-34B and further finetuned on a synthetic process reward modeling dataset (Wu et al., 2024a). We then aggregate the solutions generated by REBASE via majority voting. As shown in Figure 7, augmenting our model with REBASE scales better than majority voting, and even sequential scaling in this scenario. However, REBASE requires an additional forward pass at each step for the reward model adding some computation overhead. For sequential scaling, when prompted to use up to 512 steps, for 12 out of the 30 evaluation questions the model generates a response that exceeds the context window leading to a large performance drop. Overall, we find that these parallel scaling methods complement sequential scaling thus they offer an avenue for scaling test-time compute even further; beyond fixed context windows.

### v2

@@ 1 Introduction
Performance improvements of language models (LMs) over the past years have largely relied on scaling up train-time compute using large-scale self-supervised pretraining (Kaplan et al., 2020; Hoffmann et al., 2022). The creation of these powerful models has set the stage for a new scaling paradigm built on top of them: test-time scaling. The aim of this approach is to increase the compute at test time to get better results. There has been much work exploring this idea (Snell et al., 2024; Welleck et al., 2024), and the viability of this paradigm was recently validated by OpenAI o1 (OpenAI, 2024). o1 has demonstrated strong reasoning performance with consistent gains from scaling test-time compute. OpenAI describes their approach as using large-scale reinforcement learning (RL) implying the use of sizable amounts of data (OpenAI, 2024). This has led to various attempts to replicate their models relying on techniques like Monte Carlo Tree Search (Gao et al., 2024b; Zhang et al., 2024a), multi-agent approaches (Qin et al., 2024), and others (Wang et al., 2024a; Huang et al., 2024b, 2025). Among these approaches, DeepSeek R1 (DeepSeek-AI et al., 2025) has successfully replicated o1-level performance, also employing reinforcement learning via millions of samples and multiple training stages. However, despite the large number of o1 replication attempts, none have openly replicated a clear test-time scaling behavior. Thus, we ask: what is the simplest approach to achieve both test-time scaling and strong reasoning performance?
@@
In summary, our contributions are: We develop simple methods for creating a sample-efficient reasoning dataset (§2) and test-time scaling (§3); Based on these we build s1-32B which is competitive with o1-preview (§4); We ablate subtleties of data (§5.1) and test-time scaling (§5.2). We end with a discussion to motivate future work on simple reasoning (§6). Our code, model, and data are open-source at https://github.com/simplescaling/s1.
@@ 3.2 Metrics
Figure 4: Sequential and parallel test-time scaling. (a): Budget forcing shows clear scaling trends and extrapolates to some extent. For the three rightmost dots, we prevent the model from stopping its thinking 2/4/6 times, each time appending “Wait” to its current reasoning trace. (b): For Qwen2.5-32B-Instruct we perform 64 evaluations for each sample with a temperature of 1 and visualize the performance when majority voting across 2, 4, 8, 16, 32, and 64 of these.
@@ 4.1 Setup
We benchmark s1-32B against: OpenAI o1 series (OpenAI, 2024), which are closed-source models that popularized the idea of test-time scaling; DeepSeek r1 series (Team, 2024a), which are open-weight reasoning models with up to o1-level performance, concurrently released to ours; Qwen’s QwQ-32B-preview (Team, 2024b), a 32B open-weight reasoning model without disclosed methodology; Sky-T1-32B-Preview (Team, 2025) and Bespoke-32B (Labs, 2025), which are open models with open reasoning data distilled from QwQ-32B-preview and r1; Google Gemini 2.0 Flash Thinking Experimental (Google, 2024), the API that we distill from. As it has no official evaluation scores, we use the Gemini API to benchmark it ourselves. However, the “recitation error” of the Gemini API makes evaluation challenging. [^2] https://github.com/google/generative-ai-docs/issues/257 We circumvent this, by manually inserting all 30 AIME24 questions in its web interface where the error does not appear. However, we leave out MATH500 (500 questions) and GPQA Diamond (198 questions), thus they are N.A. in LABEL:tab:perf. Our model, s1-32B, is fully open including weights, reasoning data, and code.
@@
Model # ex. AIME 2024 MATH 500 GPQA Diamond
@@ 4.2 Performance
In Figure 2 (right) and LABEL:tab:perf we compare s1-32B with other models. We find that s1-32B is the most sample-efficient open data reasoning model. It performs significantly better than our base model (Qwen2.5-32B-Instruct) despite just training it on an additional 1,000 samples. The concurrently released r1-32B shows stronger performance than s1-32B while also only using SFT (DeepSeek-AI et al., 2025). However, it is trained on 800 $\times$ more reasoning samples. It is an open question whether one can achieve their performance with just 1,000 samples. Finally, our model nearly matches Gemini 2.0 Thinking on AIME24. As s1-32B is distilled from Gemini 2.0, this shows our distillation procedure was likely effective.
@@
[AIME24] Alice and Bob play the following game. A stack of $n$ tokens lies before them. The players take turns with Alice going first. On each turn, the player removes either $1$ token or $4$ tokens from the stack. Whoever removes the last token wins. Find the number of positive integers $n$ less than or equal to $2024$ for which there exists a strategy for Bob that guarantees that Bob will win the game regardless of Alice’s play. The problem asks for the number of positive integers $n\leq 2024$ such that Bob has a winning strategy. This is a typical impartial game that can be analyzed using the Sprague-Grundy theorem or by directly finding the winning and losing positions. Let $W$ be the set of winning positions (number of tokens $n$ such that the player whose turn it is can force a win) and $L$ be the set of losing positions. […] Let $P(n)$ be the statement ”the player whose turn it is when there are $n$ tokens has a winning strategy. $P(0)$ is false (the player whose turn it is and faces 00 tokens has already lost). Consider the small values of $n$: $n=1$: The player whose turn it is can take $1$ token and win. So $P(1)$ is true. $1\in W$. $n=2$: The player whose turn it is can take $1$ token, leaving $1$ token for the opponent. Since $1\in W$, the opponent wins. So $P(2)$ is false. $2\in L$. $n=3$: The player whose turn it is can take $1$ token, leaving $2$ tokens for the opponent. Since $2\in L$, the player whose turn it is wins. So $P(3)$ is true. $3\in W$.[…] We are looking for the number of positive integers $n\leq 2024$ for which Bob has a winning strategy. This means that if the game starts with $n$ tokens, and Alice goes first, Bob can win. This happens if and only if $n$ is a losing position (because if $n\in L$, then Alice’s first move must lead to a winning position for Bob). […] Let’s try to find a pattern in the losing positions $L=\{0,2,5,7,10,12,15,17,20,22,...\}$. Differences between consecutive terms: $2,3,2,3,2,3,2,3,2,...$ The pattern of differences is alternating $2,3$. […] The losing positions are of the form $5m$ or $25m+$ 2, where $m\geq 0$. We need to find the number of positive integers $n\leq 2024$ that are in the set of losing positions $L$. […] Case 1: $n=5m$, where $m\geq 1$. $1\leq 5m\leq 2024$ $404.81/5\leq m\leq 2024/5=$ 404.8 Since $m$ is an integer, $1\leq m\leq 404$. The number of values of $m$ is $404$. Case 2: $n=5m+2$, where $m\geq 0$. $5m+2\leq 2024$ $5m\leq 2022$ $m\leq 2022/5=404.4$ Since $m$ is an integer, $0\leq m\leq 404$. The number of values of $m$ is $404-0+1=405$. The total number of positive integers $n\leq 2024$ that are losing positions is $404+405=\boxed{809}$. [MATH500] Find the projection of $a$ onto $b=\begin{pmatrix}2\\ 6\\ 3\end{pmatrix}$ if $a\cdot b=8.$ The projection of vector $\mathbf{a}$ onto vector $\mathbf{b}$ is given by the formula: proj_b a = abb2 b We are given $\mathbf{b}=\begin{pmatrix}2\\ 6\\ 3\end{pmatrix}$ and $\mathbf{a}\cdot\mathbf{b}=8.$ First, we need to find the magnitude squared of $\mathbf{b}$, which is $\|\mathbf{b}\|^{2}$. ∥b∥^2 = bb = 2^2 + 6^2 + 3^2 = 4 + 36 + 9 = 49. Now we can plug the values into the projection formula: proj_b a = 849 b Substitute the vector $\mathbf{b}$: proj_b a = 849 (2 6 3 ) Multiply the scalar by each component of the vector: proj_b a = (849×2 849×6 849×3 ) = (164948492449) Final Answer: The final answer is ($164948492449$ ) [GPQA] A quantum mechanical particle of mass $m$ moves in two dimensions in the following potential, as a function of the polar coordinates $(r,\theta)$: V(r, θ) = 12 k r^2 + 32 k r^2 cos^2(θ) Find the energy spectrum. Hint: Write the potential in Cartesian coordinates. The problem asks for the energy spectrum of a particle in a given 2D potential. The potential is given in polar coordinates $(r,\theta)$: V(r, θ) = 12 kr^2 + 32 kr^2 cos^2(θ). We can rewrite the potential as: V(r, θ) = 12 kr^2 (1 + 3cos^2(θ)). The Hamiltonian for a particle in 2D is given by: H = -ℏ22m ( ∂2∂x2 + ∂2∂y2 ) + V(x, y). In polar coordinates $(r,\theta)$, the Laplacian operator is: ∇^2 = 1r ∂∂r (r ∂∂r ) + 1r22θ2. So the Hamiltonian in polar coordinates is: H = -ℏ22m [ 1r ∂∂r (r ∂∂r ) + 1r22θ2 ] + V(r, θ). The time-independent Schrödinger equation is $H\psi=E\psi$. We can try to separate variables by assuming a solution of the form $\psi(r,\theta)=R(r)\Theta(\theta)$. […] This confirms the energy spectrum is: E = ℏkm (2n_x + n_y + 32). Final Answer: $E=\hbar\sqrt{\frac{k}{m}}\left(2n_{x}+n_{y}+\frac{3}{2}\right)$.
@@
Figure 5: Example model outputs. We pick a question from AIME24 (left), MATH500 (middle), and GPQA (right), where our model generates the correct answer. The black text is the prompt, the light blue text is the reasoning trace, and the blue text is the answer of s1-32B. The gray ellipsis […] indicates that the text was trimmed to fit this page, but the generated text is actually longer.
@@ 5.1 Data Quantity, Diversity, and Difficulty
Table 2: s1K data ablations. We budget force (BF) a maximum of around 30,000 thinking tokens for all scores in this table. This performs slightly better than the scores without BF (LABEL:tab:perf) as it allows the model to finish with a best guess when stuck in an infinite loop. We report 95% paired bootstrap confidence intervals for differences relative to the s1K model using 10,000 bootstrap samples. E.g., the interval [-13%, 20%] means that, with 95% confidence, the true difference between 59K-full and s1K is between -13% and +20%. If the entire interval is negative, e.g. [-27%, -3%], we can confidently say that the performance is worse than s1K.
Model AIME 2024 MATH 500 GPQA Diamond
1K-random 36.7 90.6 52.0
[-26.7%, -3.3%] [-4.8%, 0.0%] [-12.6%, 2.5%]
1K-diverse 26.7 91.2 54.6
[-40.0%, -10.0%] [-4.0%, 0.2%] [-10.1%, 5.1%]
1K-longest 33.3 90.4 59.6
[-36.7%, 0.0%] [-5.0%, -0.2%] [-5.1%, 10.1%]
59K-full 53.3 92.8 58.1
[-13.3%, 20.0%] [-2.6%, 2.2%] [-6.6%, 8.6%]
s1K 50.0 93.0 57.6
@@ 5.2 Test-time scaling methods
Model AIME 2024 MATH 500 GPQA Diamond
@@ 6.1 Sample-efficient reasoning
To evaluate and push the limits of these models, increasingly challenging benchmarks have been introduced, such as Olympiad-level science competitions (He et al., 2024; Jain et al., 2024; Zhong et al., 2023) and others (Srivastava et al., 2023; Glazer et al., 2024; Su et al., 2024; Kim et al., 2024; Phan et al., 2025). To enhance models’ performance on reasoning-related tasks, researchers have pursued several strategies: Prior works have explored continuing training language models on specialized corpora related to mathematics and science (Azerbayev et al., 2023; Yang et al., 2024), sometimes even synthetically generated data (Yu et al., 2024). Others have developed training methodologies specifically aimed at reasoning performance (Zelikman et al., 2022, 2024; Luo et al., 2025; Yuan et al., 2025; Wu et al., 2024a). Another significant line of work focuses on prompting-based methods to elicit and improve reasoning abilities, including methods like Chain-of-Thought prompting (Wei et al., 2023; Yao et al., 2023a, b; Bi et al., 2023; Fu et al., 2023; Zhang et al., 2024b; Xiang et al., 2025; Hu et al., 2024). These combined efforts aim to advance the reasoning ability of language models, enabling them to handle more complex and abstract tasks effectively.
@@ Methods
As we introduce in §3, we differentiate two methods to scale test-time compute: parallel and sequential. The former relies on multiple solution attempts generated in parallel and selecting the best outcome via specific criteria. These criteria include choosing the most frequent response for majority voting or the best response based on an external reward for Best-of-N (Brown et al., 2024; Irvine et al., 2023; Snell et al., 2024). Unlike repeated sampling, previous sequential scaling methods let the model generate solution attempts sequentially based on previous attempts, allowing it to refine each attempt based on previous outcomes (Snell et al., 2024; Hou et al., 2025; Lee et al., 2025). Tree-based search methods (Gandhi et al., 2024; Wu et al., 2024b) offer a hybrid approach between sequential and parallel scaling, such as Monte-Carlo Tree Search (MCTS) (Liu et al., 2024; Zhang et al., 2023; Zhou et al., 2024; Choi et al., 2023) and guided beam search (Xie et al., 2023). REBASE (Wu et al., 2024b) employs a process reward model to balance exploitation and pruning during tree search. Empirically, REBASE has been shown to outperform sampling-based methods and MCTS (Wu et al., 2024b). Reward models (Lightman et al., 2023; Wang et al., 2024b, c) play a key role in these methods. They come in two variants: outcome reward models and process reward models. Outcome reward models (Xin et al., 2024; Ankner et al., 2024) assign a score to complete solutions and are particularly useful in Best-of-N selection, while process reward models (Lightman et al., 2023; Wang et al., 2024b; Wu et al., 2024b) assess individual reasoning steps and are effective in guiding tree-based search methods.
@@
Parallel scaling offers one solution to the limits of sequential scaling, thus we augment our sequentially scaled model with two methods: (I) Majority voting: After generating $k$ solutions, the final solution is the most frequent one across generations; (II) Tree search via REBASE: We use the REBASE process reward model, which is initialized from LLaMA-34B and further finetuned on a synthetic process reward modeling dataset (Wu et al., 2024b). We then aggregate the solutions generated by REBASE via majority voting. As shown in Figure 7, augmenting our model with REBASE scales better than majority voting, and even sequential scaling in this scenario. However, REBASE requires an additional forward pass at each step for the reward model adding some computation overhead. For sequential scaling, when prompted to use up to 512 steps, for 12 out of the 30 evaluation questions the model generates a response that exceeds the context window leading to a large performance drop. Overall, we find that these parallel scaling methods complement sequential scaling thus they offer an avenue for scaling test-time compute even further; beyond fixed context windows.